OpenStudy (anonymous):

x over 2x+1 - 2 = -4 over 7(x-2) solve

OpenStudy (mertsj):

\[\frac{x}{2x+1}-2=\frac{-4}{7(x-2)}\]

OpenStudy (mertsj):

\[\frac{x-2(2x+1)}{2x+1}=\frac{-4}{7(x-2)}\]

OpenStudy (mertsj):

\[\frac{x-4x-2}{2x+1}=\frac{-4}{7(x-2)}\]

OpenStudy (mertsj):

\[\frac{-3x-2}{2x+1}=\frac{-4}{7x-14}\]

OpenStudy (mertsj):

\[(-3x-2)(7x-14)=-8x-4\]

OpenStudy (mertsj):

\[-21x^2+28x+28=-8x-4\] \[-21x^2+36x+32=0\] \[x=\frac{6}{7}\pm\frac{2\sqrt{249}}{21}\]